OpenStudy (khally92):

Find the equation of the tangent to y=x2-3x-4 is parrelto the line y=7x+3..... i jus need to know if the answer i got is correct. i have y=7x-29 points on the curve arex=5 and y=6 are the answers correct. good evening

OpenStudy (anonymous):

You are correct.

OpenStudy (khally92):

serious

OpenStudy (khally92):

i need to differenciate this y/3x-2 i get stuck

OpenStudy (anonymous):

yes you are correct. Here is your graph. The purple line is 7x-29. The blue line is 7x+3 and the red is x^2-3x-4

OpenStudy (khally92):

i need to differenciate this y=2/3x-2... the first one was type wrong

OpenStudy (anonymous):

is it y= (2/3x) - 2 or is it y = 2/(3x-2) ?

OpenStudy (khally92):

the last one

OpenStudy (anonymous):

Differentiate using the product rule / div rule if you had (a/b) then it would be ( (a')(b) - (a)(b') )/ (b^2)

OpenStudy (anonymous):

get it ?

OpenStudy (khally92):

NOPE

OpenStudy (anonymous):

try plugging in like the example I did let your a = 2 and your b = 3x-2 a' = derivative of a b' = derivative of b b^2 = (3x-2)^2

OpenStudy (khally92):

are u aware that (3x-2) is now carryn a negative sign

OpenStudy (anonymous):

the solution is negative yes. the derivative of (3x-2) is not though. tell me the full answer u got.

OpenStudy (khally92):

i got -3(3x-2)^-2-(3x-2)^-1

OpenStudy (khally92):

-6((3x-2)^-2-(3x-2)^-1

OpenStudy (anonymous):

let me help u a bit. You were kind of on the right track. The derivative of a constant though is zero. So your setup should be something like this a = 2, derivative of a or a' = (0) b = (3x-2), derivative of b or b' = (3) so you have ( (a')(b) - (a)(b') )/(b^2) ( (0)(3x-2) - (2)(3) ) / (3x-2)^2 you should be able to take it from there.

OpenStudy (khally92):

thats -6/(3x-2)^2

OpenStudy (anonymous):

You got it. That's the answer.

OpenStudy (khally92):

find the points on the curve y=2/3x-2 where the tangent is parrellel tothe line y=-3/2x-1.....

OpenStudy (khally92):

first of ol i need to defferenciate both equations and equate them ryt?ith the ans of the first derivative is making it looks funny

OpenStudy (khally92):

Hello u still der?

OpenStudy (anonymous):

for your second equation is that -(3/2)x - 1?

OpenStudy (khally92):

yeah the derivative been -3/2

OpenStudy (anonymous):

well then after that you want to find the point where the slop is -3/2 so you set that one equal to your other derivative then solve to get your x values. Plugs back into original question to get your y values.

OpenStudy (anonymous):

* once you get your x values plug them back into your original equation to get your y values.

OpenStudy (khally92):

its not working break down the simplifyn for me here is the equation -(3/2)=-6/(3x-2)^2

OpenStudy (anonymous):

cross multiply first. if you have (a/b) = (c/d) then you cross mutliply making it a*b = b*c

OpenStudy (anonymous):